OpenStudy (katecc379):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x − y = −9 4x − 3y = −22

OpenStudy (welshfella):

8x - y = -9 we need to isolate y so first subtract 8x from both sides 8x - 8x - y = -9 - 8x - y = -8x - 9 now to get y we divide both sides of the above by -1

OpenStudy (welshfella):

just divide -y, -8x and -9 by -1

OpenStudy (katecc379):

so y=8x+9

OpenStudy (welshfella):

correct

OpenStudy (katecc379):

4(−8x − 9) − 3y = −22 4x − 3(−8x − 9) = −22 4(8x + 9) − 3y = −22 4x − 3(8x + 9) = −22 those are the only options

OpenStudy (katecc379):

so confused

OpenStudy (katecc379):

@welshfella

OpenStudy (welshfella):

ok well those are the equations after you've plugged in y = 8x + 9

OpenStudy (welshfella):

- you plug it into the second equation

OpenStudy (welshfella):